there are n bags such that i’th  bag contains i black and 2 white balls.Two balls are drawn from a randomly selected bag out of n bags.fInd the probability that both balls drawn are white.
 
Answer =1/(n+2)
 
i’m getting a different answer.

Grade:12th pass

2 Answers

Praveen Kumar Das
3 years ago
Ihere are n bags such that i’th  bag contains i black and 2 white balls.Two balls are drawn from a randomly selected bag out of n bags.fInd the probability that both balls drawn are whiteGOT THE ANSWER 1/(N+2) NOW...Thanks
Ronak Singhvi
3 years ago
Assume that i have selected i’th bag 
probability of selecting 1 bag from  n bags =1/n
probability of selecting 2 white balls at a time from the selected bag = 2C2/i+2C = 2/(i+2)(i+1)  =[1/(i+1) – 1/(i+2) ]×2
toytal probability= 2/n ∑ (1/(i+1) – 1/(i+2) ] = 2/n [ ½ – 1/n+2] = 1/n+2